# Model-free policy gradient for discrete-time mean-field control

### v1

@@ 5.2.1 Proofs on Estimation Biases
$\leq\sum_{t=0}^{T-1}\left(\left\|\mathbb{E}\left[\nabla_{l}\log\psi(\theta,X_{t+1}^{\theta},t,X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta})G_{t}^{\theta}\right]\right\|_{\infty}+\left\|\mathbb{E}\left[\nabla_{l}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta})\right]\right\|_{\infty}\right)E_{t}(\theta,\varepsilon)$
$\quad+\left\|\mathbb{E}\left[\nabla_{l}\tilde{g}(X_{T}^{\theta},l_{T}^{\theta})\right]\right\|_{\infty}E_{T}(\theta,\varepsilon)+(A_{3}+A_{5,\varepsilon}+A_{6,\varepsilon}+A_{7,\varepsilon})\varepsilon+A_{4,\varepsilon}\varepsilon^{2}$
$\leq\sum_{t=0}^{T-1}\left(B_{1}d|\mathcal{A}|(T-t+1)M_{0}+M_{1}\right)E_{t}(\theta,\varepsilon)+M_{1}E_{T}(\theta,\varepsilon)+\tilde{A}_{1,\varepsilon}\varepsilon+\tilde{A}_{2,\varepsilon}\varepsilon^{2},$
@@
$k_{1,\varepsilon}$ $\coloneqq\kappa_{1}(K_{1,\varepsilon}+K_{2,\varepsilon})\frac{2\beta_{\varepsilon}-1}{(\beta_{\varepsilon}-1)^{2}}+\kappa_{2}\beta_{\varepsilon}\frac{K_{1,\varepsilon}+K_{2,\varepsilon}}{\beta_{\varepsilon}-1},$ (5.15)
$k_{2,\varepsilon}$ $\coloneqq 2\kappa_{1}K_{1,\varepsilon}+\kappa_{2}K_{1,\varepsilon},$
$k_{3,\varepsilon}$ $\coloneqq-\frac{\kappa_{1}}{2}K_{2,\varepsilon}T^{2}-\frac{(3\kappa_{1}+2\kappa_{2})}{2}K_{2,\varepsilon}T-\kappa_{2}K_{2,\varepsilon}+\tilde{A}_{1,\varepsilon},$
$k_{4,\varepsilon}$ $\coloneqq\tilde{A}_{2,\varepsilon},$
@@ 5.2.2 Proofs on Estimation Errors
$Q_{s,i,j}^{\theta,\varepsilon,t}\coloneqq\varepsilon^{-1}\mathbf{1}_{Y_{t}^{\theta,\varepsilon,t}=x^{(i)}}\left(\Lambda_{s}^{t}\widehat{\nabla_{\theta}l^{\theta}_{s}}\right)_{j},\quad S_{s,i,j}^{\theta,\varepsilon,t}\coloneqq\mathbf{1}_{Y_{t}^{\theta,\varepsilon,t}=x^{(i)}}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta,\varepsilon,t},Y_{s}^{\theta,\varepsilon,t},l_{s}^{\theta}+\varepsilon\Lambda_{s}^{t})_{j},$
@@
$\hskip-14.22636pt\leq\frac{1}{\varepsilon^{2}\mathbb{P}\left(X_{t}^{\theta}=x^{(i)}\right)^{2}}\left(\sum_{s=0}^{t-1}{\mathrm{Var}}\left(\sum_{i^{\prime}=1}^{d}{\mathbb{E}}\left[\mathbf{1}_{Y_{t}^{\theta,\varepsilon,k}=x^{(i)}}\left(\Lambda_{s}^{t}\right)_{i^{\prime}}\right]\left(\widehat{\nabla_{\theta}l_{s}^{\theta}}\right)_{i^{\prime},j}\right)^{1/2}\right)^{2}.$
By writing $v^{\theta,\varepsilon}_{s,i^{\prime},j}={\mathrm{Var}}((\widehat{\nabla_{\theta}l_{s}^{\theta}})_{i^{\prime},j})$, we have
@@
$c_{1}$ $\coloneqq\frac{dT}{\varepsilon^{2}}\sup_{0\leq t\leq T}\sum_{i=1}^{d}\mathbb{P}\left(X_{t}^{\theta}=x^{(i)}\right)^{-2}>0,$ (5.18)
$c_{6}$ $\coloneqq\frac{1}{n\varepsilon^{2}}\sup_{0\leq t\leq T}\mathbb{E}\left[\left(\sum_{s=0}^{t-1}\sum_{i^{\prime}=1}^{d}\left|\widehat{\nabla_{\theta}l_{s}^{\theta}}\right|_{i^{\prime},j}+\varepsilon C\right)^{2}\right]\sum_{i=1}^{d}\mathbb{P}\left(X_{t}^{\theta}=x^{(i)}\right)^{-2}\in(0,\infty).$
@@
$\tilde{Q}_{t,j}^{\theta,\varepsilon}\coloneqq\varepsilon^{-1}G_{t}^{\theta,\varepsilon}\left(\Lambda_{t}\widehat{\nabla_{\theta}l^{\theta}_{t}}\right)_{j},\quad\tilde{S}_{t,j}^{\theta,\varepsilon}\coloneqq\mathbf{1}_{t<T}G_{t}^{\theta,\varepsilon}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta,\varepsilon},t,Y_{t}^{\theta,\varepsilon},l_{t}^{\theta}+\varepsilon\Lambda_{t})_{j},$
@@
$\mathcal{C}_{t}$ $=\binom{T+1}{t+1}\left(\gamma_{2}T\right)^{t}=\binom{T+1}{t+1}\left(dT\sup_{0\leq t\leq T}\sum_{i=1}^{d}\mathbb{P}\left(X_{t}^{\theta}=x^{(i)}\right)^{-2}\right)^{t}.$

### v2

@@ 5.2.1 Proofs on Estimation Biases
$\qquad\leq\sum_{t=0}^{T-1}\left(\left\|\mathbb{E}\left[\nabla_{l}\log\psi(\theta,X_{t+1}^{\theta},t,X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta})G_{t}^{\theta}\right]\right\|_{\infty}+\left\|\mathbb{E}\left[\nabla_{l}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta})\right]\right\|_{\infty}\right)E_{t}(\theta,\varepsilon)$
$\qquad\quad+\left\|\mathbb{E}\left[\nabla_{l}\tilde{g}(X_{T}^{\theta},l_{T}^{\theta})\right]\right\|_{\infty}E_{T}(\theta,\varepsilon)+(A_{3}+A_{5,\varepsilon}+A_{6,\varepsilon}+A_{7,\varepsilon})\varepsilon+A_{4,\varepsilon}\varepsilon^{2}$
$\qquad\leq\sum_{t=0}^{T-1}\left(B_{1}d|\mathcal{A}|(T-t+1)M_{0}+M_{1}\right)E_{t}(\theta,\varepsilon)+M_{1}E_{T}(\theta,\varepsilon)+\tilde{A}_{1,\varepsilon}\varepsilon+\tilde{A}_{2,\varepsilon}\varepsilon^{2},$
@@
$k_{1,\varepsilon}$ $\coloneq\kappa_{1}(K_{1,\varepsilon}+K_{2,\varepsilon})\frac{2\beta_{\varepsilon}-1}{(\beta_{\varepsilon}-1)^{2}}+\kappa_{2}\beta_{\varepsilon}\frac{K_{1,\varepsilon}+K_{2,\varepsilon}}{\beta_{\varepsilon}-1},$ (5.15)
$k_{2,\varepsilon}$ $\coloneq 2\kappa_{1}K_{1,\varepsilon}+\kappa_{2}K_{1,\varepsilon},$
$k_{3,\varepsilon}$ $\coloneq-\frac{\kappa_{1}}{2}K_{2,\varepsilon}T^{2}-\frac{(3\kappa_{1}+2\kappa_{2})}{2}K_{2,\varepsilon}T-\kappa_{2}K_{2,\varepsilon}+\tilde{A}_{1,\varepsilon},$
$k_{4,\varepsilon}$ $\coloneq\tilde{A}_{2,\varepsilon},$
@@ 5.2.2 Proofs on Estimation Errors
$Q_{s,i,j}^{\theta,\varepsilon,t}\coloneq\varepsilon^{-1}\mathbf{1}_{Y_{t}^{\theta,\varepsilon,t}=x^{(i)}}\left(\Lambda_{s}^{t}\widehat{\nabla_{\theta}l^{\theta}_{s}}\right)_{j},\quad S_{s,i,j}^{\theta,\varepsilon,t}\coloneq\mathbf{1}_{Y_{t}^{\theta,\varepsilon,t}=x^{(i)}}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta,\varepsilon,t},Y_{s}^{\theta,\varepsilon,t},l_{s}^{\theta}+\varepsilon\Lambda_{s}^{t})_{j},$
@@
$\qquad\leq\frac{1}{\varepsilon^{2}\mathbb{P}\left(X_{t}^{\theta}=x^{(i)}\right)^{2}}\left(\sum_{s=0}^{t-1}{\mathrm{Var}}\left(\sum_{i^{\prime}=1}^{d}{\mathbb{E}}\left[\mathbf{1}_{Y_{t}^{\theta,\varepsilon,k}=x^{(i)}}\left(\Lambda_{s}^{t}\right)_{i^{\prime}}\right]\left(\widehat{\nabla_{\theta}l_{s}^{\theta}}\right)_{i^{\prime},j}\right)^{1/2}\right)^{2}.$
By writing $v^{\theta,\varepsilon}_{s,i^{\prime},j}=(\Sigma_{s}^{\theta,\varepsilon})_{i^{\prime},j}={\mathrm{Var}}((\widehat{\nabla_{\theta}l_{s}^{\theta}})_{i^{\prime},j})$, we have
@@
$c_{1}$ $\coloneq\frac{dT}{\varepsilon^{2}}\sup_{0\leq t\leq T}\sum_{i=1}^{d}\mathbb{P}\left(X_{t}^{\theta}=x^{(i)}\right)^{-2}>0,$ (5.18)
$c_{6}$ $\coloneq\frac{1}{n\varepsilon^{2}}\sup_{0\leq t\leq T}\mathbb{E}\left[\left(\sum_{s=0}^{t-1}\sum_{i^{\prime}=1}^{d}\left|\widehat{\nabla_{\theta}l_{s}^{\theta}}\right|_{i^{\prime},j}+\varepsilon C\right)^{2}\right]\sum_{i=1}^{d}\mathbb{P}\left(X_{t}^{\theta}=x^{(i)}\right)^{-2}\in(0,\infty).$
@@
$\tilde{Q}_{t,j}^{\theta,\varepsilon}\coloneq\varepsilon^{-1}G_{t}^{\theta,\varepsilon}\left(\Lambda_{t}\widehat{\nabla_{\theta}l^{\theta}_{t}}\right)_{j},\quad\tilde{S}_{t,j}^{\theta,\varepsilon}\coloneq\mathbf{1}_{t<T}G_{t}^{\theta,\varepsilon}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta,\varepsilon},t,Y_{t}^{\theta,\varepsilon},l_{t}^{\theta}+\varepsilon\Lambda_{t})_{j},$
@@
$\mathcal{C}_{t}$ $=\binom{T+1}{t+1}\left(\gamma_{2}T\right)^{t}=\binom{T+1}{t+1}\left(dT\sup_{0\leq t\leq T}\sum_{i=1}^{d}\mathbb{P}\left(X_{t}^{\theta}=x^{(i)}\right)^{-2}\right)^{t},$
which concludes the proof. ∎